If 5 samples from 5 populations are given, then what is the probability of getting no Type 1 error?




1
Expert's answer
2021-06-08T18:37:52-0400

Solution:

Given data is insufficient.

If "\alpha" is level of significance, then "\alpha" is the probability of making type I error.

So, the probability of making no type I error "=1-\alpha"